Graph the piecewise function below

h(x) = { y= 3x + 2 if -2 < x < 1

y= 1/2x - 1 if x < -2

What is domain? (Interval notation)

What is range? (Interval notation)


1
Expert's answer
2021-09-19T18:19:46-0400

"h(x) = \begin{cases}
 y= 3x + 2 &\text{if } -2 < x < 1 \\
 y= \frac{1}{2}x - 1 &\text{if } x < -2
\end{cases}\\"


Now, domain for the given function is "x\in" "(-\infty,-2)\cup(-2,1)" [From graph]

At "x=1," we get "y=5"

Range is "(-\infty,-4)\cup(-4,-2)\cup(-2,5)"